### v2

# FedQUIT: On-Device Federated Unlearning via a Quasi-Competent Virtual Teacher

Abstract
Federated Learning (FL) systems enable the collaborative training of machine learning models without requiring centralized collection of individual data. FL participants should have the ability to exercise their right to be forgotten, ensuring their past contributions can be removed from the learned model upon request. In this paper, we propose FedQUIT, a novel algorithm that uses knowledge distillation to scrub the contribution of the data to forget from an FL global model while preserving its generalization ability. FedQUIT directly works on client devices that request to leave the federation, and leverages a teacher-student framework. The FL global model acts as the teacher, and the local model works as the student. To induce forgetting, FedQUIT tailors the teacher’s output on local data (the data to forget) penalizing the prediction score of the true class. Unlike previous work, our method does not require hardly viable assumptions for cross-device settings, such as storing historical updates of participants or requiring access to proxy datasets. Experimental results on various datasets and model architectures demonstrate that (i) FedQUIT outperforms state-of-the-art competitors in forgetting data, (ii) has the exact computational requirements as a regular FedAvg round, and (iii) reduces the cumulative communication costs by up to 117.6$\times$ compared to retraining from scratch to restore the initial generalization performance after unlearning.
1 Introduction
In 2016, the European Union introduced the General Data Protection Regulation (GDPR), incorporating guidelines designed to empower individuals with full control over their personal data [7]. One of the cornerstones of GDPR is the principle of data minimization, which involves collecting and processing only the personal data that is necessary for the intended purpose and limiting the retention of such data to the minimum necessary time period. Such a requirement is not easily verifiable once the data are outsourced from the data owners’ machines or devices. Especially in the case of datacenter training of Machine Learning (ML) or Deep Learning (DL) models, the collected data may be easily reused to train ML/DL models with scopes different from what was initially declared.
Federated Learning (FL) [24], an alternative to traditional centralized ML, allows the training of models while minimizing the personal information disclosed by the users. In a nutshell, FL learns a shared global model by periodically gathering and aggregating ephemeral model updates locally computed on private data, thus avoiding the transfer of unprocessed data.
The right to be forgotten is another pivotal requirement described in the GDPR. Specifically, Article 17 of GDPR states that the data subject shall have the right to obtain from the controller the erasure of personal data where the data subject withdraws consent on which the processing is based. However, when data are used to train DL models, simply deleting the samples for which the permission has been revoked is not enough. DL models can memorize and then leak private information contained in the training corpus, as widely demonstrated during the last years, e.g., via membership inference attacks (MIAs) [28, 30, 29]. Under this premise, Machine Unlearning (MU) [8, 34] has gained traction as a set of mechanisms to selectively remove the contribution of specific training samples on a learned model without destroying the generalization ability of such model. The straightforward solution would be to retrain the model, excluding the data that must be forgotten. Still, this approach can be extremely costly and sometimes unfeasible, particularly when the training dataset is no longer available. While in centralized machine learning, this naive retraining strategy might be a viable, though inefficient, option, it becomes impractical in FL. This impracticality arises due to clients’ sporadic participation, especially in cross-device federated settings [2, 16]. Consequently, developing unlearning mechanisms specifically designed for federated environments is crucial. Additionally, unlike MU, Federated Unlearning (FU) must consider that the so-called retain data, i.e., the data that remains in the process, is not directly accessible by design [27].
No Historical No Full No Stateful No Proxy No Comp. Work Information Participation Clients Data Intensive Liu et al. [22] $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ FedEraser [20] $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ FUKD [32] $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ PGA [12] $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ FedQUIT (ours) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 1: Comparison with other work performing client unlearning in federated settings.
In this paper, we present an unlearning method tailored for FL settings. Unlike recent works (e.g., [4, 20, 32, 12]), our solution is lightweight and practical for any federated client wishing to exit the process by removing its contribution. It requires no historical update storage, stateful clients, full client participation, or proxy data.
Contributions. The contributions of the paper are summarized as follows:
  • We propose FedQUIT, a novel method for FU. FedQUIT balances forgetting and knowledge retaining using a specially crafted version of the current FL global model that locally works as virtual teachers for the unlearned model in a Knowledge Distillation (KD) framework.
  • FedQUIT produces unlearned models with superior forgetting effectiveness compared to the state-of-the-art, achieving a smaller gap in forgetting performance compared to the gold-standard retrained model.
  • FedQUIT can quickly recover the gold-standard generalization ability and significantly reduce the communication cost compared to a state-of-the-art baseline. FedQUIT achieves up to 117.6$\times$ communication cost reduction compared to retraining from scratch.
  • We conduct extensive experiments on image classification tasks to demonstrate the superior forgetting capability of our method.
2 Related Work
FU approaches and their limits.
Table 1 provides a comparison with existing client unlearning methods, emphasizing key design limitations. An extended explanation of the column definitions can be found in Appendix 10. The method in [22] enables rapid retraining from scratch using historical information and Quasi-Newton methods but incurs significant computational overhead. Moreover, note that: (1) linking per-client update histories to specific clients requesting unlearning undermines FL’s privacy design; (2) to some extent it violates the ephemeral updates requirement of the FL framework [16]; and (3) at scale, considering a massive number of clients [6], several FL rounds, and multiple simultaneous learning tasks, the corresponding storage requirements might become impractical. FedEraser [20], a pioneering client unlearning approach, recalibrates historical updates but still requires the server to store all past client updates. FUKD [32] involves two server-side steps: first, it removes the unlearned client’s updates from the global model to create a sanitized model; then, it uses the previous global model as a teacher in a KD framework to improve the sanitized model’s performance, employing unlabeled proxy data. However, this approach assumes access to unlabeled data on the server, which may not always be feasible, and again requires storing historical updates. The approach in [12] resembles ours in terms of the protocol, where unlearning is performed locally by the client. Halimi et al. apply projected gradient ascent (PGA) to eliminate the influence of the data to be forgotten, constraining local weight updates to ensure that the model weights remain within an $l2$-norm ball around a reference model. However, PGA requires clients to store their latest model updates (stateful clients) and requires full client participation, as the unlearning process begins with a reference model derived by removing the last unlearning client’s model update from the current global model. In addition, PGA needs to tune, at least, a threshold that measures the distance between the unlearned model and the reference model to early stop the unlearning. Other works such as FAST [11] and the work in [1] primarily focus on backdoor removal, and, respectively, use historical information and suppose to have the unlearning client dataset divided in benign and trigger subsets. Unlike previous methods, our solution does not require stateful clients, full participation, or access to proxy data on the server. Furthermore, our method does not need to store the entire history of per-client model updates and link them to the client that requests unlearning.
Related work in centralized MU. The most similar work to ours in non-federated settings is the mechanism presented in [5], where the authors use an incompetent teacher on forget data and a competent teacher, i.e., the model before unlearning, on retain data. However, this approach is not directly applicable in federated settings since retain data – clients’ data – are not accessible. For completeness, in the experimental results, we adapt our solution to centralized settings and compare it to [5], showing superior forget effectiveness and robustness to further fine tuning after forget.
3 Background
3.1 Federated Learning
Considering a federation of $K$ clients, with each client $k$ holding a private dataset $D_{k}$, FL aims at minimizing a global objective function $f(w)$. $f(w)$ is expressed as a weighted average over all clients’ local datasets, as follows:
$\min_{w}f(w):=\sum_{k=1}^{K}\frac{n_{k}}{n}F_{k}(w),$ (1)
where $w$ represents the parameters of the global model to be trained, $n_{k}$ represents the size of $D_{k}$, and $n$ is the total number of examples held by all participating clients. $F_{k}$ is a local objective function. Federated Averaging (FedAvg) [24] is a synchronous client-server FL algorithm that operates in rounds of specific steps to solve the optimization problem of Eq. 1. At each round, the global model is distributed to a random subset of clients, which in turn locally fine-tune the model for $E$ epochs. Then, the activated clients send back a model update. The parameter server collects and aggregates the updates using weighted averaging based on the amount of per-client data samples. Finally, the server applies the averaged updates to the current global model [26].
3.2 Federated Unlearning
If FL clients are allowed to request unlearning, a given client $u$ can submit an unlearning request at any round. The unlearning client $u$ may want to have the contribution of a subset $S_{u}$ of its private data $D_{u}$, termed forgetting dataset, scrubbed from the global model. To respond to the unlearning request, an unlearning algorithm $\mathcal{U}$ is executed on the global model at a certain round $t$. Hence, $\mathcal{U}$ is applied to the global model $w_{t}$, which has been trained including client $u$ in the federation of learners, producing a scrubbed global model $w^{\bar{u}}_{t}$, i.e. $w^{\bar{u}}_{t}=\mathcal{U}(w_{t})$. After the application of $\mathcal{U}(\cdot)$, the unlearned model $w^{\bar{u}}_{t}$ should exhibit performances indistinguishable or approximately indistinguishable from a global model $w^{r}_{t}$, termed retrained model, learned without the contribution of $S_{u}$.
FU can have different objectives, i.e., sample unlearning, class unlearning [31], feature unlearning [9], and client unlearning. In this paper, we will consider the client unlearning case, where $S_{u}=D_{u}$. In practice, a client $u$ requests that the contribution of its entire local dataset is removed from the global model. Similarly to regular MU, the union of the remaining clients’ dataset $D_{r}$ is termed as retain dataset, i.e., $D_{r}=\bigcup_{\forall k\neq u}D_{k}$. The problem of FU is to design an effective and efficient forgetting mechanism to produce $w^{\bar{u}}_{t}$ from $w_{t}$. While the efficiency of an FU method is usually measured as the amount of FL rounds needed to reach the generalization performance of the retrained model (e.g., the test accuracy or loss), the ways to assess forgetting efficacy are more varied, but usually compares the unlearned model’s performances on the forgetting data with the retrained model’s ones [27].
4 FedQUIT: Unlearning via a Quasi-Competent Virtual Teacher
We design a set of FU methods, FedQUIT, that runs directly on the unlearning client’s device and fully adheres to FL privacy requirements. FedQUIT leverages modified outputs from the current global model as a teacher to achieve a balance between forgetting the client’s local data and preserving generalization on the retain data.
In this section, we begin by introducing preliminary concepts and notation (Section 4.1). Next, we detail the proposed methodology (Section 4.2). Finally, we present our set of methods, FedQUIT-Logit and FedQUIT-Softmax, which offer alternative approaches for crafting the global model output to guide the unlearning process.
4.1 Preliminaries
[…]
Figure 1: FedQUIT overview. (1) Regular training via FedAvg is performed. (2) A client $u$ submits an unlearning request to the server. (3) The server initiates a special round, in which only client $u$ participates, and runs FedQUIT locally starting from the current global model. Client $u$ sends back the unlearned model $w^{\bar{u}}$. (4) Regular training resumes from the unlearned model $w^{\bar{u}}$.
4.2 Local Unlearning
Figure 1 illustrates the unlearning process via FedQUIT. The client that requested the removal of its contribution performs the unlearning routine itself during a special round, before the regular FL training resumes. During that special round, client $u$ pulls down the last version of the global model at the time of the unlearning request, performs FedQUIT locally, and pushes the unlearned model back to the server. Then, the server can resume the regular FL training starting from the received sanitized model.
This entails that the unlearning method has direct access only to the target client’s data (the forget data), while the rest of the data in the federation (the retain data) would not be available, except via some form of proxy. Access to the knowledge derived from retain data during the unlearning phase is crucial to preserve the generalization performance of the global model while selectively removing the contribution of the forget data. A natural proxy for the data in the federation is represented by the current global model, and in our proposed method, we exploit it. As we discuss in Appendix 9, previous work has used the global model as a guide locally at clients, for example, to preserve knowledge in the presence of catastrophic forgetting and heterogeneous data. Our novel family of unlearning algorithms, FedQUIT, leverages a KD-based [14] teacher-student framework, where the global model works as teacher and the local model works as student. To induce forgetting, FedQUIT tailors the output of the teacher on forget data – the local data – penalizing the prediction score of the true class (with various possible methods explored below in this Section). In this sense, we refer to such a crafted version of the global model, as quasi-competent virtual teacher. It is worth noting that, in standard FL algorithms such as FedAvg [24], the global model is always sent to the selected clients as the starting point for the training round, so actively using it for inference on local data does not imply any further assumption over the regular FL procedure. An extended discussion on the inspirations for the design of FedQUIT can be found in Appendix 8.
Algorithm 1 Local unlearning routine in FedQUIT.
1:Current global model $w$, client (forget) data $D_{u}$, number of unlearning epochs $E_{u}$, learning rate $\eta_{u}$, batch size B
2:Unlearned model $w^{\bar{u}}$
3:Client $u$ receives the global model $w$ from the server
4:$w^{\bar{u}}\leftarrow w$triangleright Client $u$ initializes unlearned model
5:for each local unlearning epoch $e_{u}=1,2,\dots,E_{u}$ do
6:     for each minibatch =$\mathcal{B}=\{(x_{i},y_{i})\}_{i=1}^{B}\in D_{u}$ do
7:         $\hat{y}\leftarrow$ Predict($w$, $x$)
8:         $\tilde{y}\leftarrow$ ModifyOutputs($\hat{y}$, $y$)
9:         $\bar{y}\leftarrow$ Predict($w^{\bar{u}}$, $x$)
10:         Compute the loss $\mathcal{L}=\mathrm{KL}(\tilde{y}||\bar{y})$ as in Eq. 2
11:         Compute $\nabla_{w^{\bar{u}}}\mathcal{L}$
12:         $w^{\bar{u}}\leftarrow$ ClientOpt$(w^{\bar{u}},\nabla_{w^{\bar{u}}}\mathcal{L},\eta_{u})$
13:Client $u$ sends the model $w^{\bar{u}}$ back to the server
Algorithm 1 outlines the unlearning routine. To enhance readability, we have omitted the subscript $t$, which indicates the current round, from the symbols. It is worth noting that, at the beginning of the unlearning round, the unlearned model is initialized with the global model weights (line 2), i.e., the unlearned model and the global model have the same model weights when the unlearning phase starts. As anticipated, FedQUIT performs local training within a KD framework during the unlearning phase. To distill knowledge between the virtual teacher and the student, we compute their outputs on local data (respectively, line 5 and 7). The virtual teacher takes the output $\hat{y}=g(x)$ of the global model in round $t$ and alters it to produce a modified output probability $\tilde{y}=g^{\prime}(x)$ (line 6). Then, the unlearned model (the student) mimics the outputs of the virtual teacher on forgetting data $x$ (local data) by employing the following loss function:
$\mathcal{L}=\mathrm{KL}(g^{\prime}(x)||z(x))$ (2)
where $\mathrm{KL}$ is the Kullback-Leibler divergence (KL divergence), $\bar{y}=z(x)$ indicates the unlearned model’s output probability on local sample $x$.
We propose two alternative unlearning mechanisms to construct the modified probability $g^{\prime}(x)$ in the routine ModifyOutput of Alg. 1: FedQUIT-Logits and FedQUIT-Softmax. The first one works directly on logits, while the latter works on the probability output after the softmax layer.
We focus on an FL setting with possible resource-constrained client devices. Modifying the last layer of the network incurs negligible computational overhead, as it only requires adjusting a single value at the end of the forward pass (logits or softmax). Both approaches, FedQUIT-Logit and FedQUIT-Softmax, locally leverage KD to erase client contributions and only use the client data $u$’ and the current version of the global model to perform unlearning. No historical information is required, nor are proxy/public datasets supposed to exist. Next, we explain our original approaches in full detail.
FedQUIT-Logits.
Indicating (for ease of notation) the global model at round $t$ as $w_{t}$ and its output probability as $g_{t}(x)$, we design a modified output probability $g^{\prime}_{t}(x)$. Without loss of generality, we will omit the $t$ index to simplify notation. FedQUIT-Logits sets to a fixed value of $v$ the true-class logit computed by the global model before applying the softmax activation, i.e., the modified logits $z_{c}^{g^{\prime}}$ to compute the teacher’s output probabilities $g^{\prime}(x)$ become:
$z_{c}^{g^{\prime}}=\begin{dcases}v,&\text{if }c=y\\ z_{c}^{g},&\text{if }c\neq y\end{dcases}$ (3)
$\forall c\in\{1,...,C\}$. Then, $g^{\prime}(x)$ is regularly computed via softmax on $z_{c}^{g^{\prime}}$. Figure 2 provides a schematic visualization of FedQUIT-Logits.
The $v$ value is a tuning parameter that penalizes the true-class output. We empirically explored the sensitivity of FedQUIT-Logits to several choices of $v$, and in Section 6.1, we report the most robust ones that provide the best results.
FedQUIT-Softmax. FedQUIT-Softmax uses the same KD-based framework of FedQUIT-Logits, with the global model at round $t$ that works as the teacher and the unlearned model that works as the student. Differently from FedQUIT-Logit, FedQUIT-Softmax modifies the output of the global model (the teacher) after the softmax has been applied to its logits. As before, we indicate the output probabilities of the global model at round $t$ with $g(t)$. FedQUIT-Softmax replaces the true-class output probability $g(y)$ with a fixed value of $v$ and uniformly distributes the remaining quantity $g(y)-v$ to all the other class probabilities $g(c)$.
$g^{\prime}(c)=\begin{dcases}v,&\text{if }c=y\\ g(c)+\frac{g(y)-v}{C-1},&\text{if }c\neq y\end{dcases}$ (4)
The $v$ value plays the same role as in FedQUIT-Logits but takes different values that we empirically explored in Section 6.1.
Figure 2: FedQUIT-Logits.
Incompetent Teacher. As a baseline for FedQUIT-Logits and FedQUIT-Softmax, we also implemented a naive solution that uses an incompetent teacher. This alternative does not use the global model at round $t$ to craft the teacher’s output probabilities for unlearning samples, but it always outputs equally distributed probabilities over classes as an incompetent teacher, i.e., $g^{\prime}(c)=1/C$. This baseline will help us to investigate whether using the modified global model’s output as a teacher is beneficial. An incompetent teacher can also be implemented by using an untrained model, which produces random predictions.
5 Empirical Results
5.1 Settings, datasets and baselines
We conduct experiments on federated versions of two datasets, CIFAR-10 and CIFAR-100 [18]. To simulate a federated environment, we partition each dataset among 10 clients, ensuring no repeated samples. We consider both IID and non-IID data distributions. In the non-IID setting, we introduce data heterogeneity using distribution-based label skew, controlled via Latent Dirichlet Allocation [15], with concentration parameters of $\alpha=0.3$ for CIFAR-10 and $\alpha=0.1$ for CIFAR-100. Fig. 6 in the Appendix illustrates the label distribution across clients.
We employ a standard ResNet-18 [13] and a visual transformer, namely the MiT-B0 [33]. Unless stated otherwise, before performing unlearning, we train ResNet-18 from scratch for 200 FedAvg rounds while we fine-tune MiT-B0 for 50 FedAvg rounds, starting from a pre-trained checkpoint, with one local training epoch per round ($E=1$). Additionally, we validate our results using a higher number of local epochs ($E=10$) for the CIFAR-100/ResNet-18 setting with non-IID data ($\alpha=0.1$). Table 2 outlines the various settings of our experiments. For each setting, we conduct 10 experiments, each with a different client designated as the unlearning client, and we report the mean and standard deviation of the results.
Model Dataset Clients Non-IIDness Pre-trained Local Epochs ResNet-18 CIFAR-10 10 LDA ($\alpha$=0.3) No 1 ResNet-18 CIFAR-100 10 LDA ($\alpha$=0.1) or IID No 1 or 10 MiT-B0 CIFAR-100 10 LDA ($\alpha$=0.1) Yes 1
Table 2: Federated settings of reported results.
Setting Algorithm Rounds ($\downarrow$) CE ($\uparrow$) Test Acc. Forget Acc. MIA [29] MIA [36] CIFAR-100, IID, $E$=1 Original 59.9±0.0 79.8±0.7 78.2±0.6 71.4 ±0.6 Retrain 58.3±0.5 58.0±0.7 55.6±0.7 48.8 ±0.6 PGA [12] 12.1 ±4.2 16.7$\times$ 59.1±0.6 72.8 (14.8 ±1.0) 71.8 (16.1±1.8) 63.0 (14.1±1.6) Incompetent T. 15.6 ±3.8 12.8$\times$ 58.6 ±0.2 68.8 (10.8 ±2.0) 67.6 (12.0 ±1.8) 52.2 (13.4 ±1.8) FedQUIT 10.0±1.4 20.0$\times$ 58.7 ±0.4 59.4 (1.4±0.8) 60.7 (5.1±1.6) 55.7 (6.8±0.8) CIFAR-100, Non-IID, $E$=1 Original 53.8±0.0 62.9±6.9 75.4±7.0 61.0 ±7.8 Retrain 51.0±1.4 33.5±4.5 44.0±5.5 32.0 ±5.2 PGA [12] 7.0±5.1 28.6$\times$ 51.4 ±0.6 37.4 (4.3±4.1) 49.2 (6.4±4.9) 36.6 (5.3±5.3) Incompetent T. 1.7±1.25 117.6$\times$ 52.05±0.82 48.41 (14.94±6.05) 61.03 (17.04±6.50) 45.39 (13.38±6.02) FedQUIT 3.7±2.9 54.1$\times$ 52.0±0.5 34.6 (3.5±2.4) 50.7 (5.5±2.8) 33.9 (3.8±2.8) CIFAR-100, Non-IID, $E$=10 Original 50.3±0.0 60.8±6.6 73.4±6.6 61.3 ±7.5 Retrain 48.0±0.8 31.6±4.5 42.1±5.2 31.2 ±4.9 PGA [12] 4.4±3.8 45.5$\times$ 49.0±1.0 37.3 (6.2±8.1) 49.6 (7.7±9.6) 37.7 (6.6±8.2) Incompetent T. 1.0 ±0.0 200.0$\times$ 49.6 ±0.5 48.0 (16.6 ±6.4) 60.6 (18.4 ±7.7) 45.7 (14.6 ±6.7) FedQUIT 1.7±0.5 117.6$\times$ 48.3 ±0.4 27.9 (3.4±2.4) 45.9 (4.7±2.8) 29.9 (4.1±2.9) CIFAR-10, Non-IID, $E$=1 Original 83.72 ±0.00 88.59 ±3.85 84.39 ±5.35 81.37 ±6.28 Retrained 83.49 ±1.57 81.14 ±8.00 73.12 ±13.68 72.49 ±10.78 PGA [12] 11.5±5.6 17.4$\times$ 83.99 ±1.02 84.33 (3.19 ±4.86) 78.50 (5.38 ±7.70) 77.01 (4.51 ±6.04) Incompetent T. 1.1 ±0.3 181.8$\times$ 84.78 ±0.94 88.18 (7.04 ±5.08) 81.36 (8.24 ±5.96) 81.98 (9.49 ±6.34) FedQUIT 3.8 ±1.9 52.6$\times$ 83.74 ±1.17 81.71 (2.43 ±2.69) 72.12 (3.45 ±2.43) 71.06 (2.82 ±1.50)
Table 3: Performance of FedQUIT and other baselines across different settings, including data distribution and the number of local epochs, with ResNet-18. Communication Efficiency (CE column). Values in parentheses indicate the average absolute difference from the retrained model. Notably, the baselines’ Test Acc. is, by definition, higher than that of the retrained model, as we consider the recovery phase complete once test accuracy surpasses that of the retrained model.
We compare our method with the Incompetent Teacher baseline (Section 2) and the state-of-the-art PGA baseline [12]. To the best of our knowledge, PGA is the only approach that avoids impractical assumptions (see Table 2 for an extended comparison), such as requiring access to the full history of client updates, which necessitates linking clients to their past updates and compromises the privacy design of federated learning. For fairness, we enforce full client participation, as required by PGA, even though our method naturally supports the standard partial participation setting. For FedQUIT, we always use one single epoch for unlearning ($E_{u}=1$) at the target client. We report the results for FedQUIT implemented as FedQUIT-Logits with $v=0$, as it performed best. In Sec. 6.1, we also discuss the results of FedQUIT-Softmax and other choices of $v$ value. See Appendix 8 for hyper-parameter tuning details.
5.2 Metrics
We evaluate the considered methods from two perspectives: effectiveness to forget and efficiency to recover performance.
It should be noted that for all the metrics considered here to evaluate the efficacy of unlearning, a more favorable value of these metrics for an approximate unlearning method (such as ours) should minimize the absolute difference with the gold-standard retrained model.
Accuracy on forgetting data. We monitor the unlearned model’s accuracy on client $u$’s train data, and we track the gap with the retrained model [21]. This metric provides insight into how closely the unlearning process approximates the desired outcome achieved through retraining.
Membership Inference Attacks (MIA). To assess unlearning, we also use the MIA success rate. MIAs attempt to determine whether an individual sample was included in the training set of a model. The success rate of these attacks indicates how many data samples in $D_{u}$ can be correctly predicted as training records of $w_{u}$. We consider as most effective the approach that minimizes the absolute difference with the retrained model’s MIA value. We consider two different MIAs: (1) a confidence-based MIA predictor [29] and (2) a loss-based MIA predictor [36]. Appendix 12 provides detailed description of the attacks.
Recovery rounds. FU mechanisms need a recovery period (recovery phase) to match the performance of the retrained model [27]. For the federated settings, we monitor how many communication rounds the unlearned model needs to match or surpass the accuracy on the test data (referred to as Test Acc.) of the retrained model. Note that the reported Test Acc. for FU baselines is higher by design, as recovery ends once it exceeds the retrained model’s accuracy.
Communication Efficiency (CE). Communication efficiency is computed as communication rounds needed to retrain the model over recovery rounds.
5.3 Main Results
Figure 3: Test accuracy degradation after unlearning a client’s data from the FL global model, before recovery begins. Lower values indicate better preservation of generalization ability.
(a) CIFAR-100 (IID), ResNet-18.
(b) CIFAR-100 (Non-IID), ResNet-18.
(c) CIFAR-100 (Non-IID, $E$ = 10), ResNet-18.
Figure 4: Visualization of FedQUIT and PGA performance across settings. A smaller polygon indicates better unlearning effectiveness.
Algorithm Rounds ($\downarrow$) CE ($\uparrow$) Test Acc. Forget Acc. Original 75.03 ±0.0 84.25 ±5.6 Retrained 73.30 ±0.8 57.80 ±5.3 PGA [12] 6.9 ±4.3 7.2$\times$ 73.60 ±0.4 62.15 (4.35 ±3.3) FedQUIT 8.7 ±2.5 5.7$\times$ 73.42 ±0.9 58.50 (1.86 ±1.5)
Table 4: Performance of FedQUIT and other baselines on CIFAR-100 (Non-IID) using MiT-B0 as model architecture.
First, we evaluate FedQUIT in federated settings, comparing it with two baselines: the Incompetent Teacher baseline (Section 2) and the state-of-the-art PGA baseline [12]. Table 3 presents results on federated CIFAR-100 and CIFAR-10 using ResNet-18 as the model architecture, across different levels of data heterogeneity and local epochs. Table 4 reports results for federated CIFAR-100 with a visual transformer (extended results in Appendix Table 9). In Appendix Table 11, we also report results for the Natural baseline [27] as a reference, which corresponds to simply excluding the client that should be forgotten from the federation without applying any unlearning method, demonstrating that active unlearning is needed.
Forgetting effectiveness. As shown in Tables 3 and 4, FedQUIT closely approximates the performance of the gold-standard retrained model while significantly outperforming other baselines. Notably, under homogeneous data distribution (1st row of Table 3), FedQUIT is the only approach capable of fully forgetting the unlearned client’s contribution. Fig. 4 provides a visual comparison between FedQUIT and PGA, highlighting FedQUIT’s advantages in two key areas: faster recovery (fewer rounds required) and improved approximation of forgetting metrics (lower $\Delta$ in Forget Acc. and MIAs). Overall, FedQUIT shows the most effective forgetting across all settings.
Global model’s generalization ability. Fig. 3 shows accuracy degradation immediately after unlearning, before recovery begins. Unlike PGA and Incompetent Teacher, which severely harm the global model’s generalization, FedQUIT minimizes test accuracy loss, demonstrating more selective removal of client contributions.
Performance consistency during the recovery phase. Due to its lower initial degradation, FedQUIT maintains a more functional global model throughout the recovery phase.
Figure 5: Test Accuracy (Left) and Forget Accuracy (Right) for a representative client on CIFAR-100, ResNet-18, Non-IID, $E=1$.
Fig. 7 illustrates a representative recovery process starting at round 0, showing the global model’s test and forget accuracy over recovery rounds. While PGA’s model loses generalization for multiple rounds before recovering, FedQUIT experiences only minor degradation, preserving its usability for inference even during the recovery phase. Similar trends are observed across all settings (see Appendix Figs 9, 8).
Computational and communication efficiency. All federated results presented consider FedQUIT using a single local unlearning epoch ($E_{u}=1$) — the minimal possible setting. Notably, even when standard FedAvg training uses a higher number of local epochs ($E=10$), as shown in the 3rd row of Table 3, FedQUIT remains effective with just one unlearning epoch.
A similar observation applies to per-round communication costs: FedQUIT’s special unlearning round, like the other baselines, incurs the same cost as a regular FedAvg round. This is because the original model is sent to the unlearned client, and the updated global model (or its update) is sent back to the server after local unlearning. FedQUIT incurs no additional computational or communication overhead beyond standard FedAvg rounds.
While FedQUIT’s communication cost for the unlearning round is the same as FedAvg, FedQUIT significantly reduces cumulative communication costs over retraining, as shown in the CE column of Table 3. In fact, it achieves 20$\times$ reduction with IID data and up to 117.6$\times$ reduction with non-IID data. FedQUIT also outperforms PGA in this regard, with the only exception being the case MiT-B0/CIFAR-100 (see Table 4 and Appendix Table 9). In this setting, PGA shows a slightly higher gain but fails to match the retrained model’s forgetting metrics, with a higher delta in Forget Acc. indicating a less effective forgetting.
6 FedQUIT Analysis
In this section, we first examine FedQUIT’s sensitivity to the $v$ value, followed by an evaluation of its forget robustness during fine-tuning.
6.1 Sensitivity analysis on $v$
We presented results for FedQUIT implemented as FedQUIT-Logits with $v=0$, as it yielded the best performance. Here, we also analyze the impact of $v$ on FedQUIT-Softmax. Table 5 reports the performance of FedQUIT-Logits and FedQUIT-Softmax across different choices of $v$ (see also the extended results in Appendix Table 10). For FedQUIT-Logits, we examine $v=0$, that deactivates the output node corresponding to the true class, removing its contribution from the probability computation, and $v=\min\limits_{c}(z^{g})$, assigning the lowest confidence score to the true class (indicated with the notation min), to confuse it with the least likely class. For FedQUIT-Softmax, we consider $v=1/C$ and $v=0$. The former represents a uniform probability distribution across all classes, while the latter assigns null probability to the true class. Interestingly, both approaches show low sensitivity to the considered $v$ values.
IID Non-IID ($\alpha$ = 0.1) Non-IID, $E$ = 10 KD Layer $v$ R. ($\downarrow$) Forget Acc. R. ($\downarrow$) Forget Acc. R. ($\downarrow$) Forget Acc. Logit 0 10.0 59.4 (1.4) 3.7 34.6 (3.5) 1.7 27.9 (3.4) Logit min 10.1 59.7 (1.7) 3.2 34.1 (3.1) 2.1 29.5 (4.1) Softmax $1/C$ 8.1 67.2 (9.3) 3.4 40.3 (6.9) 0.7 38.8 (7.4) Softmax 0 8.3 67.9 (10.0) 2.4 39.1 (6.0) 2.1 40.2 (8.8)
Table 5: FedQUIT performance for different KD layers (logits or softmax output) and values of $v$ used in Eq. 3 (FedQUIT-Logit) or Eq. 4 (FedQUIT-Softmax). R. means communication rounds.
6.2 Forget Robustness over Fine Tuning
For completeness, we evaluate the robustness of FedQUIT over additional epochs of fine tuning in a centralized setting, i.e., whether the forgetting metrics remain stable over time after unlearning or if the influence of the removed data reemerges. We use a straightforward centralized version of FedQUIT (see Sec. 13 for detail), and we compare it against the state-of-the-art MU baseline from [5] (see extended discussion in Sec. 13). We replicate the CIFAR-Super20 setup from [5], which restructures CIFAR-100’s 60,000 images into 20 super-classes, each containing five original CIFAR-100 classes (e.g., the people super-class includes baby, boy, girl, man, woman). In the experiments, we unlearn all examples of a specific sub-class (e.g., baby), making unlearning challenging since removing one sub-class may affect other visually similar samples within the same super-class. We evaluate five cases, unlearning the sub-classes baby, rocket, lamp, sea, and mushroom. Alongside the key metrics from [5], we track test accuracy on adjacent data—samples from the same super-class as the unlearned sub-class but excluding the unlearned sub-class itself. This helps assess whether the methods can selectively erase the contribution of the forgetting data without degrading related information. ResNet-18 is used as the neural network, with images resized to 224×224 and normalized.
Table 6 presents results after one epoch of fine tuning on retain data. Appendix Tables 13 - 17 provide extended results also after 2 and 3 fine-tuning epochs. While the method from [5] produces an unlearned model with retain data accuracy comparable to retrained models even before fine-tuning, it undergoes significant degradation in forget metrics when fine-tuned further on retain data. This is evident from the high positive accuracy gap on forget data after just one fine-tuning epoch, such as +34% in the lamp case. In contrast, FedQUIT preserves forget metrics near the gold standard, with only a slight increase over the second and third fine-tuning epochs, demonstrating its effectiveness and robustness in forgetting.
Task Algorithm Retain Data Adjacent Data Forget Data MIA [29] Baby Original 85.68 94.92 91.00 98.80 Retrain 85.92 ±0.21 93.64 ±0.12 82 ±0.41 73.6 ±0.22 FedQUIT (central.) 85.01 ±0.23(-0.92) 88.83 ±4.07(-4.81) 80.33 ±4.04(-1.67) 67 ±6.22(-6.6) Chundawat et al. 84.40 ±0.3(-1.53) 91.24 ±0.62(-2.4) 82 ±2(0) 77.67 ±3.21(4.07) Rocket Original 85.57 91.04 92.00 99.40 Retrain 85.39 ±0.26 81.4 ±0.32 3 ±0.34 0.8 ±0.29 FedQUIT (central.) 85.16 ±0.08(-0.23) 85.22 ±4.72(3.82) 7 ±2(4) 1.07 ±0.58(0.27) Chundawat et al. 84.90 ±0.11(-0.48) 81.94 ±1.77(0.54) 25.33 ±7.02(22.33) 6.47 ±3.8(5.67) Lamp Original 85.03 86.11 78.00 88.60 Retrain 84.93 ±0.46 90.76 ±0.31 15 ±0.18 8.2 ±0.35 FedQUIT (central.) 85.07 ±0.32(0.14) 87.64 ±4.21(-3.12) 18 ±10.54(3) 8.13 ±4.39(-0.07) Chundawat et al. 84.62 ±0.32(-0.32) 89.52 ±3.06(-1.24) 49 ±3(34) 26.93 ±4.41(18.73)
Table 6: Test accuracy on retain, adjacent, and forget data, obtained after one epoch of fine tuning on retain train data, following unlearning. The performance (%) is presented as $a\pm b(c)$, where $a$ is the mean, $b$ is the standard deviation, computed over five runs, and $c$ is the gap with the retrained model, i.e., closer is better.
7 Conclusion
We presented FedQUIT, a novel and general teacher-student framework for FU. In particular, we use crafted versions of the FL global model’s outputs as teacher to selectively perform unlearning while retaining the good knowledge of the original model. In the experiments, we show that using such specially crafted teachers is an effective approach to forget data and favors rapid recovery.
Supplementary Material
8 Inspiring Observations
We aim to design an FU method that operates on-device and fully adheres to the FL privacy requirements. This entails that the method would have direct access only to the unlearning client’s data, while the rest of the data in the federation (the retain data) would not be available for use in the unlearning algorithms, except via some form of proxy. To indirectly utilize the knowledge extracted from the retain data, we leverage KD and use the FL global model as the teacher, which serves as a proxy for the data we cannot access. However, we craft a virtual teacher by modifying the global model output, to balance forgetting of local data and retaining of global knowledge. In the following paragraphs, we first provide an overview of the usage of the global model as a local teacher in FL to combat data heterogeneity. Next, we present (and critically evaluate) the usage of an incompetent teacher to induce forgetting in centralized settings. We define our teacher as quasi-competent because its output resembles the original probability distribution for the non-true classes while penalizes the output for the true class.
Usage of competent teacher to tackle catastrophic forgetting.
In FL, and especially in cross-device FL, data are supposed to be non-IID. This characteristic translates to drift of client models, that diverge from the global model making the next FedAvg aggregation inefficient. Extensive research has investigated the use of the global model as an anchor for local models to mitigate this drift, aiming to minimize the disruptive impact of data heterogeneity on aggregation. An approach to implementing this guidance is via KD, with the global model acting as the teacher and the local model as the student [35, 19, 23, 10]. This approach preserves global knowledge while incorporating local adaptations, which helps limit the phenomenon of catastrophic forgetting – the rapid overwriting of previous knowledge to accommodate the distribution shift imposed by the specific clients participating in each round.
Usage of incompetent teacher to induce forgetting.
On the other hand, inducing forgetting – the complementary of tackling catastrophic forgetting – via KD is typically achieved using an incompetent or random teacher in place of a competent teacher. The underlying assumption of using a random or incompetent teacher is that the unlearned model should be indifferent to the forgetting data, producing non-informative predictions when presented with such data. However, this assumption often does not hold true in practice. It is common that the retrained model still produces informative predictions on forgetting data, which the retrained model have never seen during training, demonstrating its ability to generalize. This is evident from the results of our experiments in a centralized setting. In Appendix Table 8, for the rocket case, we note that the retrained model is not able to classify rockets as vehicles 2 (rockets’ super-class). This is because rockets are very diverse from the other images in the same super-class (lawn-mower, streetcar, tank, tractor). On the other hand, for the baby case (Appendix Table 5) we have that the retrained model is able to generalize on the test images of babies even though it has never seen a baby once in the training set. This is because the super-class people contains classes of images (boy, girl, man, woman) which share features with the baby images. Therefore, when considering a retrained model as the gold standard for the performance of the unlearned model, the unlearned model should not produce suspiciously random predictions on forgetting data in circumstances where it should be able to generalize instead. In our example above, even though the unlearned model should have forgotten about babies, it should still be inclined to correctly identify their super-class. Consequently, using a random or incompetent teacher is not ideal, and we investigate possibly better teachers to induce forgetting while preserving generalization performance.
8.1 Relation with Self-Supervision
In our method – and in general when the global model guides local training – at the beginning of the unlearning routine, the unlearned model and the global model (the student and the teacher) have the same model parameters. In this sense, the FedQUIT unlearning routine resembles a self-supervised learning, with the difference that the self-supervision is from a slightly modified version of self. Similarly, in FedQUIT’s adaptation to centralized settings, the original model guides itself – via modified outputs – to balance forgetting and knowledge retaining.
9 Similarly Inspired Work in FL
The mechanisms that we present in this paper use a student-teacher framework locally at FL clients to retain the good knowledge from the original model while selectively scrubbing the contributions to forget. We use a crafted version of the FL global model as the teacher, serving as a natural proxy for the retain data that cannot be directly accessed in FL. Recently, KD [14, 3] has been widely used in FL to address various issues [25], such as reducing communication costs, enabling model heterogeneity within the federation, improving generalization or personalization performance with heterogeneous data, and even performing unlearning. Most relevantly for this paper, numerous works employ the global model as a KD teacher for the local model to mitigate catastrophic forgetting and client drift in the presence of heterogeneous data (e.g., [35, 19, 23, 10]). For example, FedGKD [35] uses the global model output as a guide for local models, limiting client model drifts. FedNTD [19] demonstrated that excluding the teacher’s true-class logit during the distillation phase is beneficial for local-global distillation. Note that FedNTD [19] and FedQUIT differ significantly in both objective and methodology: FedNTD addresses data heterogeneity, while FedQUIT enables client unlearning. Technically, (1) FedNTD applies KL-divergence as a local regularizer for cross-entropy loss, whereas FedQUIT directly optimizes the KL divergence in Eq. 2; (2) FedNTD ignores the logits produced by the true classes, while FedQUIT actively penalizes them—an essential step to induce forgetting. Other recent contributions have proposed other label-masking approaches for the global model’s output (i.e., the teacher) [23, 10]. Unlike these works, our approach modifies the output of the global model, which still acts as a teacher for the local model, to perform unlearning.
10 Extended Description of Table 1
Historical information. This refers to whether a method requires storing and accessing historical data, such as the complete history of per-client updates, which is typically maintained by the parameter server. Note that: (1) Linking per-client update histories to specific clients requesting unlearning undermines FL’s privacy design; (2) to some extent it violates the ephemeral updates requirement of the FL framework [16]; and (3) at scale, considering a massive number of clients [6], several FL rounds, and multiple simultaneous learning tasks, the corresponding storage requirements might become impractical.
Stateful client. It indicates whether clients must store information, such as the last model update they have computed, after sending their contribution back. Especially in cross-device settings, clients are supposed to be stateless.
Full participation. It indicates whether the method requires that all clients in the federation participate during each round. While this is an assumption that can be made in cross-silo settings, in cross-device settings only a small subset of clients is selected per round.
Proxy data. It indicates whether the method requires access to proxy data. Proxy data should be semantically similar to the data within the federation, which can make this assumption challenging to meet.
11 Infrastructure and Libraries
We run all the experiments on a machine with Ubuntu 22.04, equipped with 64 GB of RAM and one NVIDIA RTX A5000 as GPU (32GB memory). We developed the code with Python and with Python libraries; in our code repository, we provide the instructions to exactly reproduce our Python environment.
(a) CIFAR-10 (Non-IID, $\alpha=0.3$).
(b) CIFAR-100 (Non-IID, $\alpha=0.1$).
(c) CIFAR-100 (IID).
Figure 6: Label distribution across clients (0-9) for CIFAR-10 (Non-IID, $\alpha=0.3$) and CIFAR-100 (IID and Non-IID,$\alpha=0.1$).
12 Membership Inference Attacks
In this section, we briefly describe Shokri’s attack and Yeom’s attack that we use in the experimental results. In general, MIA metrics reflect the information leakage of training algorithms about individual members of the training corpus. A lower MIA success rate implies less information about $D_{u}$ in $w_{u}$.
Song’s MIA [29].
This attack is implemented using the prediction confidence-based attack proposed in [29]. This approach includes a training phase where a balanced dataset is sampled from data seen during training and data unseen during traning. For seen data, we use the retain dataset, which includes the training data but excludes any forget data; in federated settings, this retain dataset consists of the combined data from the remaining clients. For unseen data, we use a standard test dataset that the model has never encountered during training and is distinct from the forgetting dataset. The MIA predictor is trained to identify whether the target model’s output resembles that of known (seen) data or unknown (unseen) data. To assess unlearning effectiveness, the MIA predictor is then applied to a model using the forgetting data as input.
Yeom’s MIA [36].
This attack assumes a white-box setting in which the attacker has knowledge of the model’s average training loss. A data sample is predicted to be a member of the training set if its model loss is below this average training loss; otherwise, it is classified as a non-member. It is important to note that this metric is used primarily for comparison with other baselines. In practice, an attacker would need access to each client’s local training average loss, which is generally infeasible. Alternatively, the attacker would require knowledge of the global model’s average loss across the union of all distributed client datasets, which is similarly challenging to obtain. In our implementation, we assume the latter scenario.
13 Baselines in Experiments
Natural baseline (federated baseline):
During the unlearning routine, there is no explicit unlearning. The unlearned model has the exact same model parameters of the original model. The training resumes, and the unlearning client is excluded from the federation. The rationale is that we want to understand whether an explicit unlearning algorithm is more efficient than letting the model adapt to the new data distribution (i.e., without the unlearning client).
PGA [12] (federated baseline):
The work in [12] applies projected gradient ascent (PGA) to eliminate the influence of the data to be forgotten, directly on-device (similarly to our method). PGA constrains the local weight updates to ensure that the model weights remain within an $l2$-norm ball around a carefully selected reference model. The reference model is computed by removing the last unlearning client’s model update from the current global model. Halimi et al. perform early stopping if the distance between the unlearned model and the reference model is below a threshold, which must be tuned. Note that: (1) PGA requires clients to store their last model updates to build the so-called reference model, and, to do that, it maintains a state on clients; (2) in light of the previous point, PGA needs full participation of clients; (3) PGA needs to tune two thresholds (plus the standard training hyper-parameters).
Competent-Incompetent Teacher [5] (centralized baseline):
During the unlearning process, forget data and retain data are interleaved. The model is trained to mimic the output of the original model when presented with retain data, while it emulates the output of an incompetent model (e.g., a random model) when given forget data. It is worth noting that forget data and retain data need to be properly balanced through an additional hyper-parameter that controls the proportion of the two sets. Note that FedQUIT is foundamentally different from this approach: (1) FedQUIT explicitly targets FL settings and not centralized settings, even though it can be applied in the latter setting as well (as we demonstrated in the Experimental Results Section), (2) FedQUIT does not need retain data (which are not available in FL settings when working directly on the unlearning client’s device), (3) FedQUIT tailors a virtual teacher and modifies the output of the FL global model to induce unlearning, (4) The work in [5] also needs to tune a ratio to balance the amount retain data and forget data.
Centralized version of FedQUIT.
For the experiments reported in Sec. 6.2, we use a straightforward centralized version of FedQUIT. The unlearning phase is the same of federated FedQUIT, performing unlearning via FedQUIT’s algorithm starting from the original model and using centralized forget data. The fine-tuning epochs are then performed on the resulting unlearned model, using centralized retain data.
14 Hyper-parameter Tuning and Pre-processing
In this Section, we report the hyper-parameter tuning of the various methods we used.
14.1 Regular Training (Federated Settings)
CIFAR-10/CIFAR-100 Data Distribution.
Figure 6 shows the label distribution across clients for the federated CIFAR-10 and CIFAR-100.
ResNet-18 on CIFAR-10/CIFAR-100.
We used a standard ResNet-18 [13] and employed Group Normalization layers, similar to other works with similar settings (e.g., [17]). We used SGD as a local optimizer with a learning rate set to 0.1, with a round-wise exponential decay of 0.998, 1 local epoch, local batch size of 32. We pre-processed the training images with random crop, horizontal flip and normalization layers. During unlearning routines, we only apply normalization. As server-side optimizer, we consider SGD with a learning rate set to 1.0 (standard FedAvg aggregation and update application).
MiT-B0 on CIFAR-100.
We used a visual transformer, i.e., MiT-B0 [33], with approximately 3.6M parameters, initialized from a pre-trained model checkpoint trained on ImageNet-1k (69.27% accuracy on test data). We adapted the one-layer classification head to this task, initializing such a layer from scratch. We employed the AdamW optimizer with a client learning rate of 3e-4, with a round-wise exponential decay of 0.998, 1 local epoch, local batch size of 32, and weight decay regularization of 1e-3. As server-side optimizer, we consider SGD with a learning rate set to 1.0 (standard FedAvg aggregation and update application). The images are resized to a resolution of 224x224.
14.2 Unlearning Phase (Federated Settings)
ResNet-18 on CIFAR-10/CIFAR-100.
FedQUIT (all variants), Incompetent Teacher: see Table 7.
Dataset Non-IIDness Optimizer $\eta_{u}$ $E_{u}$ Batch Size CIFAR-100 IID Adam 1e-3 1 32 CIFAR-100 Non-IID Adam 1e-4 1 32 CIFAR-10 Non-IID Adam 1e-4 1 32
Table 7: Hyper-parameter tuning for unlearning routing of FedQUIT (federated settings).
PGA [12]: the same hyper-parameter prescribed in the original paper, and we tuned their early stopping threshold in $[2.2,10.0]$ selecting the following as best: 9.0 (IID), 6.5 (Non-IID).
MiT-B0 on CIFAR-100.
FedQUIT (all variants), Incompetent Teacher: 1 unlearning epochs, learning rate 5e-4, AdamW optimizer, local batch size 32. PGA [12]: we report in Table 8 the best configs of hyperparameter (learning rate for local unlearning $\mu_{PGA}$, optimizer opt, unlearning epochs Eu, threshold $\tau$). We selected ($\mu_{PGA}$=3e-4, opt=AdamW, Eu=5, $\tau=7.5$) as best. We set batch size to 512 (similar to what prescribed in the original paper).
Algorithm Rounds $\Delta$ Forget Acc. $\Delta$ Test Acc. $\Delta$ MIA [29] PGA (1e-2, SGDM, 5, 6.5) 6.4±7.5 10.52±7.88 0.62±0.77 11.54±9.25 PGA (1e-2, SGDM 5, 7.0) 12.3±11.09 8.98±7.19 0.56±0.8 9.62±9.01 PGA (3e-4, AdamW, 5, 7.0) 3.8±2.97 8.61±6.64 0.65±0.85 8.24±7.23 PGA (3e-4, AdamW, 5, 7.5) 6.9±4.28 4.35±3.28 0.25±0.36 4.98±3.67 PGA (3e-4, AdamW, 5, 8.0) 7.5±3.74 5.06±3.11 0.33±0.27 5.81±3.67
Table 8: Comparison among different hyperparameter configs for PGA on federated CIFAR-100 using a transformer-based (MiT-B0) architecture. The values that represent the absolute difference in performance compared to the retrained model are indicated with $\Delta$.
Centralized settings.
For the baseline from [5] we used the hyperparameters from the original paper. For FedQUIT (all variants) we used the Adam optimizer with learning rate of 5e-4 (tuned in $[1e-1,1e-2,1e-3,5e-4,1e-4,1e-5]$).
15 Further Experimental Results
Figure 7: Test Accuracy (Left) and Forget Accuracy (Right) for a representative client on CIFAR-100, ResNet-18, Non-IID, $E=1$.
Figure 8: Test Accuracy (Left) and Forget Accuracy (Right) for a representative. Setting: ResNet-18, CIFAR-100, IID, $E=1$.
Figure 9: Test Accuracy (Left) and Forget Accuracy (Right) for a representative client. Setting: miT-B0, CIFAR-100, Non-IID, $E=1$.
In this Section, we include further experimental results that are mentioned in the main paper but excluded for the sake of space.
Global model during the recovery phase (all settings). Figs. 7, 8, 9 show a representative recovery process (for a random client) tracking the global model’s test and forget accuracy, for IID and Non-IID settings.
Complete results for the setting MiT-B0/CIFAR-100. Table 9 reports the complete results for the setting MiT-B0/CIFAR-100, relative to the reduced version of the table Table 4 provided in the main paper for the sake of space.
Complete results for the sensitivity analysis of FedQUIT. Table 10 reports the complete results for the sensitivity analysis of FedQUIT’s $v$ value (extended version of Table 5).
Complete results for centralized setting. Tables 13 - 17 report the full results for our centralized experiments, which we reduced to be included in the main text as in Table 6 of the main paper.
Degradation after unlearning (centralized setting). Fig. 10 report the degradation after unlearning and before recovery for the centralized setting.
Algorithm Rounds ($\downarrow$) CR ($\uparrow$) Test Acc. Forget Acc. MIA [29] Original 75.03 ±0.00 84.25 ±5.63 77.03 ±8.57 Retrained 73.30 ±0.78 57.80 ±5.29 48.59 ±4.96 PGA [12] 6.9 ±4.28 7.2$\times$ 73.60 ±0.36 62.15 (4.35 ±3.28) 53.57 (4.98 ±3.67) FedQUIT 8.78 ±2.54 5.7$\times$ 73.42 ±0.89 58.50 (1.86 ±1.46) 48.86 (2.56 ±2.13)
Table 9: Performance of FedQUIT and other baselines across different settings. Values in parentheses indicate absolute differences from the retrained model. Dataset: CIFAR-100 (Non-IID), Model architecture: MiT-B0.
Setting KD Layer $v$ Rounds ($\downarrow$) CR ($\uparrow$) Test Acc. Forget Acc. MIA [29] MIA [36] CIFAR-100, IID, $E$=1 Logit 0 10.0±1.4 20.0$\times$ 58.70 ±0.4 59.43 (1.4±0.8) 60.71 (5.1±1.6) 55.70 (6.8±0.8) Logit min 10.1±1.73 58.59±0.58 59.66 (1.70±1.01) 61.96 (6.32±1.78) 55.92 (7.08±0.86) Softmax $1/C$ 8.1±1.66 58.61±0.40 67.23 (9.26±0.85) 66.63 (11.01±1.52) 61.72 (12.88±1.10) Softmax 0 8.3±1.34 58.57±0.43 67.95 (9.98±0.91) 66.34 (10.73±1.00) 61.98 (13.14±1.08) CIFAR-100, Non-IID, $E$=1 Logit 0 3.7±2.9 54.1$\times$ 52.00±0.5 34.60 (3.5±2.4) 50.74 (5.5±2.8) 33.93 (3.8±2.8) Logit min 3.2±1.87 51.66±1.23 34.10 (3.05±2.05) 50.10 (6.08±3.32) 33.54 (3.11±2.29) Softmax $1/C$ 3.4±3.75 51.55±1.62 40.31 (6.85±3.83) 53.86 (9.8±4.11) 38.60 (6.59±3.58) Softmax 0 2.4±1.9 51.57±1.32 39.09 (5.97±3.52) 52.31 (8.39±4.62) 36.81 (5.92±2.91) CIFAR-100, Non-IID, $E$=10 Logit 0 1.7±0.5 117.6$\times$ 48.30 ±0.4 27.93 (3.4±2.4) 45.90 (4.7±2.8) 29.90 (4.1±2.9) Logit min 2.00±1.00 48.81±0.89 29.45 (4.14±3.24) 47.33 (5.29±4.70) 31.04 (3.90±3.22) Softmax $1/C$ 0.71±1.67 48.85±0.95 38.82 (7.43±4.85) 52.14 (10.02±5.18) 37.98 (6.88±5.22) Softmax 0 2.11±0.60 49.00±0.87 40.20 (8.81±5.23) 54.01 (11.78±5.57) 39.81 (8.72±5.73)
Table 10: FedQUIT performance across different KD layers (logits or softmax output) and values of $v$ used in 3 (FedQUIT-Logit) and 4 (FedQUIT-Softmax). The table reports the number of communication rounds and forgetting accuracy across IID and Non-IID settings, including scenarios with $E$=10.
Figure 10: CIFAR20-Super.
Natural baseline (federated) results. Table 11 reports the results for the Natural baseline. After 20 rounds the unlearned client has detached the federation, its contribution is, on average, significantly noticeable (high gap in forget acc and MIAs with the retrained model).
Algorithm Round Test Acc. Forget Acc. MIA [29] MIA [36] Original 53.8±0.0 62.9±6.9 75.4±7.0 61.0 ±7.8 Retrain 51.0±1.4 33.5±4.5 44.0±5.5 32.0 ±5.2 Natural 1 53.56 (2.56±1.01) 54.28 (20.78±7.23) 67.32 (23.32±7.86) 52.48 (20.48±7.33) Natural 2 53.41 (2.41±1.63) 52.05 (18.55±7.08) 64.93 (20.93±7.70) 49.91 (17.91±7.33) Natural 4 53.29 (2.29±1.45) 50.49 (16.99±6.06) 63.11 (19.11±7.19) 48.44 (16.44±6.26) Natural 16 53.31 (2.31±1.30) 45.74 (12.24±4.84) 57.86 (13.86±5.73) 43.16 (11.16±4.52) Natural 20 53.67 (2.67±1.35) 45.37 (11.87±4.84) 56.97 (12.97±5.99) 42.18 (10.18±4.80)
Table 11: Performance of Natural baseline in the setting CIFAR-100, Non-IID, $E=1$, ResNet-18. Values in parentheses indicate the average absolute difference from the retrained model. Note that still after 20 rounds the unlearned client has detached the federation, its contribution is, on average, significantly noticeable (substantial gap in forget acc and MIAs with the retrained model).
sea Retain Data Adjacent Data Forget Data MIA
Original 85.19 91.75 95.00 97.40
Retrain 85.13 ±0(0) 85.33 ±0(0) 83 ±0(0) 63.4 ±0(0)
Incompetent Teacher 83.98 ±0.28(-1.14) 87.93 ±0.76(2.6) 84.67 ±1.15(1.67) 77.33 ±1.5(13.93)
FedQUIT-Logitzero 84.96 ±0.2(-0.17) 91.1 ±1.4(5.77) 85.67 ±3.21(2.67) 63.33 ±4.15(-0.07)
FedQUIT-Softmax 1/C 84.06 ±0.45(-1.07) 87.83 ±3.1(2.5) 85.33 ±1.15(2.33) 77.07 ±5.1(13.67)
Chundawat et al. 84.31 ±0.57(-0.82) 89.72 ±1.01(4.39) 91 ±3(8) 84.2 ±4.08(20.8)
mushroom Retain Data Adjacent Data Forget Data MIA
Original 85.80 92.93 79.00 95.80
Retrain 85.75 ±0(0) 91.34 ±0(0) 1 ±0(0) 2.2 ±0(0)
Incompetent Teacher 83.4 ±0.74(-2.35) 90.62 ±2.61(-0.72) 7.33 ±3.06(6.33) 7.4 ±1.22(5.2)
FedQUIT-Logitzero 85.07 ±0.22(-0.67) 88.87 ±2.98(-2.47) 7 ±1.73(6) 2.73 ±0.12(0.53)
FedQUIT-Softmax1/C 83.51 ±0.68(-2.23) 87.91 ±5.75(-3.43) 5 ±1(4) 4 ±0.35(1.8)
Chundawat et al. 84.29 ±0.36(-1.46) 91.21 ±3.2(-0.13) 31 ±14.73(30) 12.93 ±10.25(10.73)
Table 12: CIFAR-Super20 results when unlearning sea and mushroom sub-classes.
baby Epochs Retain Data Adjacent Data Forget Data MIA
Original 85.68 94.92 91.00 98.80
Retrain 85.92 ±0(0) 93.64 ±0(0) 82 ±0(0) 73.6 ±0(0)
Incompetent 0 4.73 ±0.7(-81.19) 0 ±0(-93.64) 0 ±0(-82) 0.07 ±0.12(-73.53)
1 83.01 ±0.57(-2.92) 84.7 ±4.55(-8.94) 71.33 ±5.86(-10.67) 69 ±5.13(-4.6)
2 84.54 ±0.4(-1.38) 92.61 ±0.89(-1.03) 84.33 ±2.08(2.33) 77.73 ±1.6(4.13)
3 84.78 ±0.34(-1.15) 94.38 ±2.21(0.74) 84 ±7(2) 82.27 ±4.09(8.67)
Logit 0 26.41 ±1.66(-59.52) 61.03 ±8.3(-32.61) 59.33 ±5.69(-22.67) 67.33 ±3.04(-6.27)
1 85.01 ±0.23(-0.92) 88.83 ±4.07(-4.81) 80.33 ±4.04(-1.67) 67 ±6.22(-6.6)
2 84.87 ±0.17(-1.05) 94 ±0.59(0.36) 84 ±2(2) 64.67 ±1.86(-8.93)
3 84.76 ±0.23(-1.17) 90.6 ±0.7(-3.05) 80.67 ±0.58(-1.33) 64.6 ±2.43(-9)
Softmax 0 5.92 ±8.67(-80.0) 14.84 ±27.71(-78.8) 1 ±0.12(-81) 0.08 ±49.17(66.07)
1 82.88 ±0.49(-3.05) 89.58 ±1(-4.07) 79.33 ±0.58(-2.67) 74.93 ±3.56(1.33)
2 85.29 ±0.41(-0.63) 92.8 ±1.45(-0.84) 87.67 ±1.53(5.67) 81.73 ±3.29(8.13)
3 85.27 ±0.12(-0.66) 90.76 ±1.59(-2.89) 83.67 ±1.15(1.67) 80.07 ±3.41(6.47)
Chundawat et al. 0 84.19 ±0.08(-1.74) 91.07 ±1.28(-2.57) 78 ±1.73(-4) 0.2 ±0.35(-73.4)
1 84.4 ±0.3(-1.53) 91.24 ±0.62(-2.4) 82 ±2(0) 77.67 ±3.21(4.07)
2 84.55 ±0.13(-1.37) 91.11 ±1.57(-2.53) 86.67 ±4.04(4.67) 85.07 ±2.64(11.47)
3 84.03 ±0.93(-1.89) 93.85 ±2.98(0.21) 87.67 ±6.43(5.67) 88.53 ±5.71(14.93)
Table 13: CIFAR-Super20 extended results when unlearning baby sub-classes. In particular, the table reports the performance immediately after unlearning (after 0 epochs) and after 1, 2, 3 epochs.
sea Epochs Retain Data Adjacent Data Forget Data MIA
Original 85.19 91.75 95.00 97.40
Retrain 85.13 ±0(0) 85.33 ±0(0) 83 ±0(0) 63.4 ±0(0)
Incompetent 0 4.81 ±0.49(-80.31) 0.33 ±0.56(-85) 0.67 ±1.15(-82.33) 66.2 ±57.33(2.8)
1 83.98 ±0.28(-1.14) 87.93 ±0.76(2.6) 84.67 ±1.15(1.67) 77.33 ±1.5(13.93)
2 85.13 ±0.12(0) 90.17 ±0.49(4.84) 90 ±2(7) 76.8 ±7.03(13.4)
3 84.65 ±0.3(-0.48) 86.94 ±0.73(1.61) 89.67 ±0.58(6.67) 82.07 ±1.55(18.67)
Logit 0 13.79 ±0.6(-71.34) 8.8 ±4.25(-76.53) 7 ±2(-76) 12.8 ±7.74(-50.6)
1 84.96 ±0.2(-0.17) 91.1 ±1.4(5.77) 85.67 ±3.21(2.67) 63.33 ±4.15(-0.07)
2 84.84 ±0.11(-0.29) 90.57 ±1.65(5.24) 86 ±1.73(3) 65.8 ±1.64(2.4)
3 84.79 ±0.26(-0.34) 88.87 ±2.19(3.54) 86 ±5.29(3) 66.2 ±7.2(2.8)
Softmax 0 6.16 ±0.81(-78.97) 0.25 ±0.43(-85.08) 0.33 ±0.58(-82.67) 0 ±0(-63.4)
1 84.06 ±0.45(-1.07) 87.83 ±3.1(2.5) 85.33 ±1.15(2.33) 77.07 ±5.1(13.67)
2 85.22 ±0.44(0.09) 89.55 ±0.27(4.22) 85.67 ±4.16(2.67) 72.6 ±2.09(9.2)
3 85.06 ±0.47(-0.07) 89.34 ±0.75(4.01) 88 ±4.58(5) 78.67 ±4.03(15.27)
Chundawat et al. 0 84.83 ±0.09(-0.3) 89.58 ±3.1(4.25) 79 ±2(-4) 0 ±0(-63.4)
1 84.31 ±0.57(-0.82) 89.72 ±1.01(4.39) 91 ±3(8) 84.2 ±4.08(20.8)
2 83.67 ±0.51(-1.45) 89.72 ±1.74(4.39) 90.33 ±1.53(7.33) 89 ±2.16(25.6)
3 84.54 ±0.26(-0.58) 87.5 ±2.02(2.17) 89.33 ±4.16(6.33) 82.33 ±6.83(18.93)
Table 14: CIFAR-Super20 extended results when unlearning sea sub-classes.
lamp Epochs Retain Data Adjacent Data Forget Data MIA
Original 85.03 86.11 78.00 88.60
Retrain 84.93 ±0(0) 90.76 ±0(0) 15 ±0(0) 8.2 ±0(0)
Incompetent 0 4.62 ±0.43(-80.32) 31.42 ±54.41(-59.34) 32.67 ±56.58(17.67) 32.53 ±55.14(24.33)
1 82.73 ±0.43(-2.21) 84.5 ±2.62(-6.26) 13.67 ±3.79(-1.33) 10.67 ±3.26(2.47)
2 85.04 ±0.2(0.11) 84.87 ±0.6(-5.89) 17 ±6.08(2) 8 ±2.23(-0.2)
3 84.93 ±0.32(0) 87.85 ±2.02(-2.91) 21.33 ±4.51(6.33) 12.13 ±0.64(3.93)
Logit 0 30.33 ±1.43(-54.61) 10.45 ±3.32(-80.31) 7.33 ±0.58(-7.67) 17.73 ±5.06(9.53)
1 85.07 ±0.32(0.14) 87.64 ±4.21(-3.12) 18 ±10.54(3) 8.13 ±4.39(-0.07)
2 85.48 ±0.27(0.54) 89.61 ±1.69(-1.15) 28.33 ±12.01(13.33) 12.6 ±4.78(4.4)
3 85.02 ±0.48(0.08) 87.63 ±2.71(-3.13) 25 ±4(10) 12.33 ±0.83(4.13)
Softmax 0 7.47 ±0.59(-77.47) 6.26 ±5.37(-84.5) 1.67 ±1.53(-13.33) 1.33 ±1.53(-6.87)
1 81.53 ±0.34(-3.41) 83.87 ±0.52(-6.89) 17 ±2(2) 11.4 ±3.54(3.2)
2 84.97 ±0.36(0.03) 88.03 ±2.69(-2.73) 21 ±5.29(6) 10.6 ±3.86(2.4)
3 85.15 ±0.3(0.22) 90.84 ±1.49(0.09) 24 ±11.14(9) 15.13 ±5.43(6.93)
Chundawat et al. 0 84.8 ±0.22(-0.13) 88.55 ±3.5(-2.21) 11.67 ±4.16(-3.33) 0.13 ±0.23(-8.07)
1 84.62 ±0.32(-0.32) 89.52 ±3.06(-1.24) 49 ±3(34) 26.93 ±4.41(18.73)
2 84.93 ±0.14(0) 90.36 ±1.04(-0.4) 50.67 ±5.13(35.67) 37.47 ±4.69(29.27)
3 84.96 ±0.31(0.03) 88.58 ±1.47(-2.18) 45.33 ±13.2(30.33) 32 ±9.82(23.8)
Table 15: CIFAR-Super20 extended results when unlearning lamp sub-class.
rocket Epochs Retain Data Adjacent Data Forget Data MIA
Original 85.57 91.04 92.00 99.40
Retrain 85.39 ±0(0) 81.4 ±0(0) 3 ±0(0) 0.8 ±0(0)
Incompetent 0 5.3 ±1.17(-80.09) 0 ±0(-81.4) 0 ±0(-3) $3$57.74(32.53)
1 84.16 ±0.68(-1.23) 86.46 ±2.01(5.06) 3 ±1.73(0) 1.67 ±0.46(0.87)
2 85.17 ±0.11(-0.22) 86.7 ±4.43(5.3) 5.67 ±3.51(2.67) 2 ±1.31(1.2)
3 84.94 ±0.39(-0.45) 87.01 ±2.81(5.61) 4.33 ±2.31(1.33) 2.33 ±1.4(1.53)
Logit 0 27.48 ±0.51(-57.9) 47.52 ±7.05(-33.88) 24 ±2(21) 17.8 ±14.08(17)
1 85.16 ±0.08(-0.23) 85.22 ±4.72(3.82) 7 ±2(4) 1.07 ±0.58(0.27)
2 84.8 ±0.2(-0.59) 83.15 ±4.49(1.74) 7.67 ±2.52(4.67) 3.07 ±1.89(2.27)
3 85.01 ±0.29(-0.38) 84.56 ±3.92(3.16) 9.33 ±1.53(6.33) 3.27 ±1.22(2.47)
Softmax 0 6.37 ±0.29(-79.02) 2.03 ±1.75(-79.37) 1 ±1(-2) 10.87 ±9.1(10.07)
1 84.15 ±0.18(-1.23) 86.19 ±0.21(4.79) 4.33 ±0.58(1.33) 1.53 ±1.79(0.73)
2 85.4 ±0.46(0.01) 83.08 ±5.32(1.68) 4.67 ±1.15(1.67) 2.07 ±1.01(1.27)
3 84.89 ±0.38(-0.5) 85.09 ±2.95(3.69) 6.67 ±0.58(3.67) 3.27 ±0.99(2.47)
Chundawat et al. 0 84.59 ±0.32(-0.8) 83.46 ±3.29(2.05) 4.67 ±5.51(1.67) 0.07 ±0.12(-0.73)
1 84.9 ±0.11(-0.48) 81.94 ±1.77(0.54) 25.33 ±7.02(22.33) 6.47 ±3.8(5.67)
2 84.78 ±0.24(-0.6) 83.04 ±7.55(1.64) 31 ±7.81(28) 13.6 ±5.57(12.8)
3 84.91 ±0.26(-0.48) 84.71 ±2.68(3.31) 24 ±5(21) 9.87 ±3.5(9.07)
Table 16: CIFAR-Super20 extended results when unlearning rocket sub-class.
mushroom Epochs Retain Data Adjacent Data Forget Data MIA
Original 85.80 92.93 79.00 95.80
Retrain 85.75 ±0(0) 91.34 ±0(0) 1 ±0(0) 2.2 ±0(0)
Incompetent 0 4.51 ±1.19(-81.24) 6.12 ±10.6(-85.22) 0.67 ±1.15(-0.33) 99.93 ±0.12(97.73)
1 83.4 ±0.74(-2.35) 90.62 ±2.61(-0.72) 7.33 ±3.06(6.33) 7.4 ±1.22(5.2)
2 85.6 ±0.18(-0.15) 91.62 ±1.96(0.28) 10.67 ±5.51(9.67) 6.33 ±1.36(4.13)
3 85.09 ±0.5(-0.65) 95.17 ±0.6(3.83) 18.67 ±2.52(17.67) 12.47 ±2.7(10.27)
Logit 0 27.28 ±2.45(-58.47) 46.49 ±11.38(-44.85) 41.33 ±4.51(40.33) 52.07 ±6.8(49.87)
1 85.07 ±0.22(-0.67) 88.87 ±2.98(-2.47) 7 ±1.73(6) 2.73 ±0.12(0.53)
2 85.05 ±0.38(-0.7) 93.11 ±1.22(1.77) 18.67 ±4.16(17.67) 6.93 ±1.33(4.73)
3 84.88 ±0.4(-0.86) 92.89 ±1.18(1.55) 13 ±4.36(12) 7.13 ±0.5(4.93)
Softmax 0 4.97 ±0.83(-80.78) 31.94 ±55.32(-59.4) 33.67 ±57.45(32.67) 66 ±57.17(63.8)
1 83.51 ±0.68(-2.23) 87.91 ±5.75(-3.43) 5 ±1(4) 4 ±0.35(1.8)
2 85.18 ±0.25(-0.57) 89.38 ±2.84(-1.96) 10 ±5.2(9) 5.13 ±2.84(2.93)
3 84.89 ±0.17(-0.86) 94.41 ±0.77(3.07) 19.33 ±4.16(18.33) 9.4 ±1.91(7.2)
Chundawat et al. 0 84.7 ±0.46(-1.04) 88.94 ±2.82(-2.4) 3.67 ±1.53(2.67) 0 ±0(-2.2)
1 84.29 ±0.36(-1.46) 91.21 ±3.2(-0.13) 31 ±14.73(30) 12.93 ±10.25(10.73)
2 84.66 ±0.11(-1.09) 90.6 ±3.51(-0.74) 31.67 ±12.1(30.67) 15.27 ±9.26(13.07)
3 84.81 ±0.27(-0.93) 91.33 ±3(-0.01) 24 ±3(23) 10.53 ±2.14(8.33)
Table 17: CIFAR-Super20 extended results when unlearning mushroom sub-class.